1.a) Technology Corporation expects to order 126,000 memory chips for inventory during the coming year, and it will use this inventory at a constant rate. Fixed ordering costs are $225 per order; the purchase price per chip is $25; and the firm’s inventory carrying costs is equal to 20 percent of the purchase price. (Assume a 360-day year.) What is the economic ordering quantity for chips?

1.b) Use the data from problem 1a. If the lead time for placing an order is 15 days, and Technology Corporation holds a safety stock equal to a 30-day supply of chips, then at what inventory level should an order be placed?
Answer #1

1a: EOQ = SQRT (2*Annual demand * ordering cost/ Carrying cost)

= SQRT (2* 126000* 225/ 20%*25)

=84187.29

rounded off to

84187

b: Reorder point = Average daily sales * lead time + Safety stock

= (126000/360) * 15 + 126000*30/360

=15750